The Arthur Company manufactures kitchen utensils. The company is currently producing well below its full capacity. The Benton Company has approached Arthur with an offer to buy 23,000 utensils at $0.80 each. Arthur sells its utensils wholesale for $0.90 each; the average cost per unit is $0.87, of which $0.12 is fixed costs. If Arthur were to accept Benton's offer, what would be the increase in Arthur's operating profits?

Multiple Choice

  • $690.

  • $1,610.

  • $1,150.

  • $2,300.
Answer #1

Answer ) $1,150

The average cost per unit is 0.87

Less: Fixed cost per unit    0.12

Variable cost per unit is 0.75

The Benton company has offered $0.80 per unit

So, Profit per unit = 0.80-0.75 = $0.05 per unit

Total profit for 23000 unit = 23000*0.05

   = $1,150

Here the Average fixed cost per unit is 0.12, which will remained constant for total production.

The company does not have incurr any additional expenses in connection to fixed cost. Only variable cost has to incurr for additional production.